OpenStudy (derpisthewordorisit):

What is the third term in the pattern with formula 8 · 4n – 1? A. 64 B. 96 C. 128 D. 512

OpenStudy (derpisthewordorisit):

@XSuikodenSoulCaliburX

OpenStudy (derpisthewordorisit):

@Nnesha

OpenStudy (derpisthewordorisit):

B?

Nnesha (nnesha):

\[\huge\rm a_n = 8 \times 4_{n-1}\] like this ?

OpenStudy (derpisthewordorisit):

its just 8x4n-1

Nnesha (nnesha):

\[\huge\rm 8 \times 4n -1\] like this

OpenStudy (derpisthewordorisit):

yes

OpenStudy (derpisthewordorisit):

so B?

Nnesha (nnesha):

how did you get b?

OpenStudy (derpisthewordorisit):

well im not quite sure i dont really understand it...

OpenStudy (derpisthewordorisit):

i could be wrong

Nnesha (nnesha):

take a screenshot of the question

OpenStudy (derpisthewordorisit):

how..?

Nnesha (nnesha):

PrntScr and the paste that in paint or if u use chrome then lightshot

OpenStudy (derpisthewordorisit):

i use chrome

Nnesha (nnesha):

okay wait

OpenStudy (derpisthewordorisit):

i clicked prntscreen and nothing..

Nnesha (nnesha):

is it 8 times 4(n-1) ?

OpenStudy (derpisthewordorisit):

yes

Nnesha (nnesha):

parentheses are really important because if it's 8 times 4n-1 n=3 so u will get 95 which is not correct

Nnesha (nnesha):

so it's \[8 \times 4(n-1)\] right ?